If 5.0 mol of N2O4 is placed in a 1.0L closed container at 150C, what will be the concentrations of N2O4 and NO2 at equilibrium? K=4.5
1
Expert's answer
2020-11-04T14:24:46-0500

Solution:

The initial concentration of N2O4:

Co(N2O4) = n(N2O4) / V(N2O4) = 5.0 mol / 1.0 L = 5.0 M

Co(N2O4) = 5.0 M


N2O4(g) = 2NO2(g)

The equlibrium constant expression for the reaction:

K = [NO2]2 / [N2O4] = 4.5


........................N2O4(g) = 2NO2(g)

I (initial)..............5.0..............--.......

C (change)..........-x.............+2x......

E (equlibrium)....5.0-x..........2x.......


K = [NO2]2 / [N2O4] = (2x)2 / (5.0-x) = 4.5

4x2 = 4.5 × (5.0-x)

4x2 = 22.5 - 4.5x

4x2 + 4.5x - 22.5 = 0

x = 1.875


[N2O4] = 5.0 - x = 5.0 - 1.875 = 3.125 M

[N2O4] = 3.125 M

[NO2] = 2x = (2 × 1.875) = 3.75 M

[NO2] = 3.75 M


Answer: at equilibrium, the concentrations of NO2 and N2O4 will be 3.75 M and 3.125 M, respectively (if K = 4.5).